OpenStudy (anonymous):

What is the range when the outlier is excluded? Data set for age received driver's license: sixteen, thirteen, eighteen, seventeen, nineteen, seventeen, sixteen, eighteen, sixteen, seventeen, and thirty-one. 5 6 seventeen eighteen @mikurout

OpenStudy (anonymous):

@CrashOnce

OpenStudy (anonymous):

range formula

OpenStudy (anonymous):

tell me

OpenStudy (anonymous):

Idk...

OpenStudy (anonymous):

range= largest value-lowest value so

OpenStudy (anonymous):

31-13=18

OpenStudy (anonymous):

the outlier of data is =19-13=6